QDx=300-2Px+4Py-2Pz-.01Y where,Px=$25 Py=$35 Pz=$20 Y(income)=20000 a) calculate the QDx b) explain briefly the relationship between i) good x and good y ii) good x and good z c) is good x normal or inferior please explain
1
Expert's answer
2016-03-09T08:36:43-0500
QDx = 300 - 2Px + 4Py - 2Pz - 0.01Y where, Px=$25 Py=$35 Pz=$20 Y(income)=20000
a) QDx = 300 - 2*25 + 4*35 - 2*20 - 0.01*20000 = 150 units.
b) the relationship between goods are:
i) good x and good y are substitutes, because the increase in price for good y will cause the increase in quantity demanded of good x.
ii) good x and good z are complements, because the increase in price for good z will cause the decrease in quantity demanded of good x.
c) good x is inferior, because with the increase in income the quantity demanded of good x will decrease.